OpenStudy (anonymous):

Jeans are on sale for 20% off the regular price. Which function would give you the total cost for purchasing a pair of jeans that costs x dollars? f(x) = 0.2x f(x) = 0.2x - x f(x) = x - 0.2x

Mehek (mehek14):

the first one because you multiply .2 by the regular price

OpenStudy (anonymous):

Well, you would multiply 0.2 x the value of the jeans, then subtract the value of the jeans by that, so what is your answer?

OpenStudy (anonymous):

@Mehek14 That is incorrect

OpenStudy (anonymous):

@countryboy212 Please don't listen to @Mehek14 . He is incorrect.

OpenStudy (anonymous):

f(x) = x - 0.2x For example, let x=10 f(x)=10-0.2(10) f(x)=10-2 The sale price will be $8.

OpenStudy (anonymous):

Multiply 0.2 x the value of the jeans, then subtract the value of the jeans by that, so what is your answer?

Mehek (mehek14):

im sorry i forgot the subtracting part

OpenStudy (anonymous):

than what is it

OpenStudy (anonymous):

Look at my comment above.

Mehek (mehek14):

it would be the last one